If the required return on this stock is 14.5 percent, what is the current share price?

Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next nine years because the firm needs to plow back its earnings to fuel growth. The company will pay a dividend of $15 per share 10 years from today and will increase the dividend by 5 percent per year thereafter. If the required return on this stock is 14.5 percent, what is the current share price?


 Current share price = _______ 
Answer #1

Here we have a stock that pays no dividends for 9 years. Once the stock begins paying dividends, it will have a constant growth rate of dividends. We can use the constant growth model at that point. It is important to remember that general constant dividend growth formula is:

Pt = [Dt × (1 + g)] / (R − g)

This means that since we will use the dividend in Year 10, we will be finding the stock price in Year 9. The dividend growth model is similar to the PVA and the PV of a perpetuity: The equation gives you the PV one period before the first payment. So, the price of the stock in Year 9 will be:

P9 = D10 / (R − g) = $15 / (0.145 − 0.05) = $157.894737

The price of the stock today is simply the PV of the stock price in the future. We simply discount the future stock price at the required return. The price of the stock today will be:

P0 = $157.894737 / 1.1459 = $46.68